OpenStudy (anonymous):

Solve the following system of equations. Explain which method you chose to solve the system. 2x + y = 2

OpenStudy (anonymous):

Need more equations haha

OpenStudy (anonymous):

x - y = 10 2x + y = 2

OpenStudy (anonymous):

Add the two equations to get 3x=12 x=4 Plug x back in, 2(4) +y=2 y=-6 x=4 y=-6 Elimination and substitution.

OpenStudy (anonymous):

elimination method add the equation eliminates y so u solve for x and then find y by plug in y in 1st equation

OpenStudy (anonymous):

edge has done the algebra

OpenStudy (anonymous):

Ok Hold on... Edge. can you elaborate on add the two equations?

OpenStudy (anonymous):

x - y = 10 2x + y = 2 adding -y + y = 0 so y disappears you get 3x + 0 = 12 x = 4

OpenStudy (anonymous):

What jimmy said

OpenStudy (anonymous):

Ok Then What?

OpenStudy (anonymous):

plug x=4 in x - y = 10 so 4 - y = 10 -y = 10 - 4 = 6 y = -6